OpenStudy (anonymous):

HELP!!! On a 250 question test and I need help with the questions I cant understand. 1: Which of the following is a solution of the compound inequality? 3x + 8 < 2 or x + 12 > 2 - x Answer x < -2 or x > 5 x < -6 and x > 5 All real numbers No solution 2: Solve |x + 6| = 12. Answer 6 -18 {6, -18} No Solution

OpenStudy (anonymous):

i can try

OpenStudy (karategirl2002):

me too

OpenStudy (anonymous):

What do you the solution is to these two? 1: Which of the following is a solution of the compound inequality? 3x + 8 < 2 or x + 12 > 2 - x Answer x < -2 or x > 5 x < -6 and x > 5 All real numbers No solution 2: Solve |x + 6| = 12. Answer 6 -18 {6, -18} No Solution

OpenStudy (karategirl2002):

6+6=12

OpenStudy (karategirl2002):

for 2

OpenStudy (anonymous):

that is not an option.

OpenStudy (karategirl2002):

what do you mean

OpenStudy (anonymous):

Im in high school. math is way harder than that

OpenStudy (karategirl2002):

oh

OpenStudy (anonymous):

dont worry about it. I know you were just trying to help and be nice

OpenStudy (karategirl2002):

:)

OpenStudy (anonymous):

I really do need help though. does any body have good ideas. im horrible in math

OpenStudy (anonymous):

x < -2 or x > 5... for the first one: I graphed it :D

OpenStudy (anonymous):

thank you and what about the next one?

OpenStudy (anonymous):

6: |x + 6| = 12 x + 6 = 12 --> take out of absolute value x = 6 --> subtract 6 from each side and solution! :D

OpenStudy (anonymous):

oh that makes soooo much sense! would you be willing to help me with a few more?

OpenStudy (anonymous):

sure! :D

OpenStudy (karategirl2002):

okay

OpenStudy (anonymous):

ok here we go 3: Solve |2x -5| > 7. Answer x > 6 or x < -1 -1 < x < 6 x > 6 and x < -1 4: Determine which ordered pairs are solutions to the inequality, y < 2x - 8. Answer (5, -10) (3, -6) (-4, 0) (10, 5)

OpenStudy (anonymous):

-1 < x < 6

OpenStudy (anonymous):

for the first one

OpenStudy (anonymous):

I will give you a medal and friend you after you help me with all the questions I have if you want.

OpenStudy (anonymous):

or fan you

OpenStudy (anonymous):

Hi. I haven't looked at all of these, but I just noticed something about one of the problems above... 2: Solve |x + 6| = 12. Answer 6 -18 {6, -18} No Solution

OpenStudy (anonymous):

ok :D the other one is (5, -10)

OpenStudy (anonymous):

yeah? @TOMWALK

OpenStudy (anonymous):

is that right @MelissaHolmes

OpenStudy (anonymous):

-18 is also a possible value for x. |x + 6| = 12 |-18+6| = 12 |-12| = 12 12 = 12 So, {6, -18} is the right answer.

OpenStudy (anonymous):

okay....

OpenStudy (anonymous):

just a sec

OpenStudy (anonymous):

Which are counter examples of the following statement? If the sum of x and y is negative, then both x and y are negative. Answer x = -2 and y = -4 x = -10 and y = 5 x = -6 and y = 0 x = 15 and y = -2

OpenStudy (anonymous):

hello?

OpenStudy (karategirl2002):

yeah

OpenStudy (anonymous):

sry.. i was away for a bit..

OpenStudy (anonymous):

x = -6 and y = 0

OpenStudy (anonymous):

is that it because you can put more than one answer into that one

OpenStudy (anonymous):

yeah :D

OpenStudy (anonymous):

ok. I thought it was cool you want to become a doctor.

OpenStudy (karategirl2002):

me too

OpenStudy (anonymous):

@karategirl2002 do you play minecraft? I think you do lol

OpenStudy (karategirl2002):

lol yeah

OpenStudy (anonymous):

I just started do you have any tips. Im going to play on my ipad mini

OpenStudy (karategirl2002):

um look out for mobs the computer one is better btw you need internet to play with friend on that start out in survivle and go to creative

OpenStudy (karategirl2002):

@Coolgirl1999

OpenStudy (anonymous):

I meant that I do play on my ipad lol yah I get what you saying and the computer version is so $$$$$$. I can afford it

OpenStudy (anonymous):

cant

OpenStudy (karategirl2002):

it is 24.95 forever so

OpenStudy (anonymous):

whats you name? @karategirl2002 Im Lydia

OpenStudy (karategirl2002):

kristina nice to meet you lydia

OpenStudy (karategirl2002):

brb

OpenStudy (anonymous):

you too

OpenStudy (karategirl2002):

back

OpenStudy (anonymous):

k. what shows do you watch on tv?

OpenStudy (anonymous):

@karategirl2002

OpenStudy (anonymous):

i see you want to go to stanfod lol

OpenStudy (anonymous):

your soo cute

OpenStudy (anonymous):

yah do you watch lab rats or kickin it

OpenStudy (anonymous):

or i didnt do it

OpenStudy (karategirl2002):

@hartnn

OpenStudy (karategirl2002):

yeah

OpenStudy (anonymous):

I love kickin it and im 14 lol

OpenStudy (karategirl2002):

im going to be 12

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

u home schooled or do u go to reg school

OpenStudy (karategirl2002):

home schooled

OpenStudy (anonymous):

me too!

OpenStudy (anonymous):

what school

OpenStudy (karategirl2002):

k12.com

OpenStudy (anonymous):

for how long? I was in that school from kindergarten to 5th grade

OpenStudy (karategirl2002):

2nd then 5th and 6th

OpenStudy (anonymous):

nice, do you plan on going to high school home schooled

OpenStudy (karategirl2002):

no im going to luis center next year

OpenStudy (anonymous):

nice is that near by you

OpenStudy (anonymous):

hey I gotta go ill see you later..... ill text you some time ttyl

OpenStudy (karategirl2002):

k and yeah